The metabolism of glucose (molar mass 180.2 g/mol) with oxygen forms CO2 and H2O and releases 678 kcal/mol of energy. C6H12O6(aq) + 6 O2(g) ⇌ 6 CO2(g) + 6 H2O(l) glucose
a. Are the bonds formed in the products stronger or weaker than the bonds broken in the reactants?
Stronger or weaker?
b. How much energy is released when 6.00 mol of glucose is metabolized?
? kcal
Report answer to appropriate number of significant figures.
1
Expert's answer
2019-05-27T06:09:15-0400

a) As the energy is released then the bonds formed in the product sare stronger than the bonds broken in the reactants

Answer: stronger

b) According to equation


"C_6H_{12} O_6 + 6O_2 = 6 CO_2 + 6 H_2O +678 kcal\/mol"

1 mole of glucose releases 678 kcal of energy

we have 6.00 mol of glucose which relese x kcal of energy

Solve the proportion


"\frac{1}{6.00}=\frac{678}{x}"

"x= 4068"

we shoud round the answer to three significant figures, i.e.


"x= 4070 kcal"

Answer: 4070 kcal